OpenStudy (lxelle):

The function f is such that f(x) = a − b cos x for 0◦ ≤ x ≤ 360◦, where a and b are positive constants. The maximum value of f(x) is 10 and the minimum value is −2. Find the values of a and b.

OpenStudy (alekos):

have you tried to work it out ?

hartnn (hartnn):

think : when can f(x) be maximum ?? when cos x is minimum ? or maximum or 0 ? then using the appropriate value of cos x you can find the mim. and max. value of f(x)

OpenStudy (lxelle):

What I did was, When cosx = -1, a-b=10 and when cosx = 1, a + b = 2. But it seems to be wrong. Help? @hartnn @alekos

hartnn (hartnn):

when cos x = -1 then a+b = 10 :)

hartnn (hartnn):

a- b(-1) = 10 a+b =10

OpenStudy (lxelle):

Yup. Why is that so?

hartnn (hartnn):

a-b =2

OpenStudy (lxelle):

Isn't it maximum value of f(x) is 10 Hence cox(x)= 1???

hartnn (hartnn):

NOTE : there is -b cos x there is a negative sign! when cos x is maximum, f(x) will be minimum!

OpenStudy (lxelle):

Does that mean that cos is negative?

OpenStudy (lxelle):

because it says b is a positive constant.

hartnn (hartnn):

b is positive hence -b will be negative now there is a negative quantity with cos x so, f(x) will not increase with cos x, it decreases with cos x hence f(x) will reach maximum, when cos x is minimum

OpenStudy (lxelle):

Okay. Let's say the question is f(x)= a+bcos(x) So, It'll be a-b = 10 and a+b =-2 right?

hartnn (hartnn):

absolutely

OpenStudy (lxelle):

Appreciate it!:)

hartnn (hartnn):

^_^